# Mingda Zhang1, Na Zhao1,2, Jianglong Qin1,2,*

### v3

@@ 4.3 Manual Evaluation Results of Legal Dispute Analysis
Indicator DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B Llama 3.1 70B Mistral 7B
Base Trad. Comp. Base Trad. Comp. Base Trad. Comp. Base Trad. Comp.
Information Quality 3.2 3.8 4.1 3.8 4.1 4.4 4.1 4.3 4.6 2.9 3.5 3.9
Understanding & 2.9 3.5 3.8 3.7 3.9 4.4 4.0 4.2 4.5 2.6 3.2 3.6
Reasoning
Expression Style 3.4 3.9 4.2 4.0 4.2 4.5 4.2 4.4 4.7 3.1 3.6 4.0
Safety 3.6 4.2 4.5 4.2 4.4 4.6 4.4 4.5 4.8 3.3 3.8 4.2
Trust 3.0 3.7 4.0 3.7 4.0 4.5 4.0 4.3 4.6 2.7 3.3 3.8
Average 3.22 3.82 4.12 3.88 4.12 4.48 4.14 4.34 4.64 2.92 3.48 3.90
The manual evaluation results presented in Table 8 further confirm the limitations of single prompt methods in legal dispute analysis tasks. Traditional configuration performs lower on all QUEST evaluation dimensions, such as DeepSeek-R1-Distill-Qwen-32B model scoring 2.9 in the "Understanding & Reasoning" dimension (compared to baseline’s 2.9, showing no improvement) and 3.7 in the "Trust" dimension (23.3% improvement from baseline). Legal expert evaluation points out that content generated by traditional configuration, while containing legal terminology and citations, presents a "superficially professional, substantively shallow" characteristic when solving actual legal problems. The limitations of single prompts emphasizing knowledge acquisition while neglecting reasoning guidance become more apparent in complex legal dispute cases. In contrast, complete configuration achieves significant improvements in all evaluation dimensions, with Llama 3.1 70B achieving the highest scores across all dimensions (Information Quality: 4.6, Understanding & Reasoning: 4.5, Expression Style: 4.7, Safety: 4.8, Trust: 4.6). Even the smallest Mistral 7B model shows substantial improvements, with Trust scores increasing from 2.7 to 3.8 (40.7% improvement), demonstrating the advantages of the three-stage prompt engineering and multidimensional knowledge graph fusion framework. The framework achieves a transition from "legal knowledge application" to "legal wisdom expression," with generated content that is not only professionally accurate but also possesses sophisticated legal reasoning and perspective consideration, enhancing legal experts’ recognition of artificial intelligence systems in legal dispute analysis.
@@ 5 Conclusion
This research proposes an enhanced framework for legal dispute analysis that integrates prompt engineering with multidimensional knowledge graphs to address fundamental limitations in legal language models. The framework establishes a three-stage hierarchical prompt structure (comprising task definition, knowledge background, and reasoning guidance components) that works synergistically with a three-layer knowledge graph architecture to form a closed-loop system for legal analysis. Experimental validation confirms the framework’s effectiveness across multiple dimensions, with the complete configuration demonstrating improvements in both automated metrics and expert evaluations. The framework achieves enhanced text generation quality, superior legal judgment capabilities (with sensitivity increasing by 9.9-13.8%, specificity by 4.8-6.7%, and precision by 10.3-15.3%), as well as higher quality legal content in terms of citation accuracy (22.4-39.7

### v4

@@ 4.3 Manual Evaluation Results of Legal Dispute Analysis
Indicator DeepSeek-R1-Distill-Qwen-32B Qwen/QwQ-32B
Baseline Traditional Complete Baseline Traditional Complete
Information Quality 3.6 3.2 4.1 4.1 3.7 4.4
Understanding & Reasoning 3.3 2.9 3.8 4.0 3.6 4.4
Expression Style & Role 3.8 3.4 4.2 4.2 3.9 4.5
Safety & Harm 4.0 3.6 4.5 4.3 4.0 4.6
Trust & Confidence 3.5 3.0 4.0 4.0 3.5 4.5
The manual evaluation results presented in Table 4 further confirm the limitations of single prompt methods in legal dispute analysis tasks. Traditional configuration performs lower on all QUEST evaluation dimensions, such as DeepSeek-R1-Distill-Qwen-32B model scoring 2.9 in the "Understanding & Reasoning" dimension (12.1% lower than baseline) and 3.0 in the "Trust & Confidence" dimension (14.3% lower than baseline), indicating that single prompts limit the depth and reliability of model legal reasoning, resulting in reduced trust from legal experts in system outputs. Legal expert evaluation points out that content generated by traditional configuration, while containing legal terminology and citations, presents a "superficially professional, substantively shallow" characteristic when solving actual legal problems, with the limitations of single prompts emphasizing knowledge acquisition while neglecting reasoning guidance becoming more apparent in complex legal dispute cases. In contrast, complete configuration achieves significant improvements in all evaluation dimensions, demonstrating the advantages of the three-stage prompt engineering and multidimensional knowledge graph fusion framework, achieving a transition from "legal knowledge application" to "legal wisdom expression," with generated content that is not only professionally accurate but also possesses sophisticated legal reasoning and perspective consideration, enhancing legal experts’ recognition of artificial intelligence systems in legal dispute analysis.
@@ 5 Conclusion
This research proposes an enhanced framework for legal dispute analysis that integrates prompt engineering with multidimensional knowledge graphs to address fundamental limitations in legal language models. The framework establishes a three-stage hierarchical prompt structure (comprising task definition, knowledge background, and reasoning guidance components) that works synergistically with a three-layer knowledge graph architecture to form a closed-loop system for legal analysis. Experimental validation confirms the framework’s effectiveness across multiple dimensions, with the complete configuration demonstrating improvements in both automated metrics and expert evaluations. The framework achieves enhanced text generation quality, superior legal judgment capabilities (with sensitivity increasing by 7.3-11.3%, specificity by 5.4-6.0%, and precision by 8.7-12.5%), as well as higher quality legal content in terms of citation accuracy, reasoning rationality, and conclusion reliability.